OpenStudy (anonymous):

multiply and simplify 10/51 3/8 this product is ?

OpenStudy (anonymous):

10/51 * 3/8 = 5/68 you can cross reduce so 10/51 becomes 5/17 and 3/8 becomes 1/4 so 5/17*1/4 = 5/68

OpenStudy (anonymous):

or you can just multiply straight across and simplify at the end...